Problem 04-3A Applying activity-based costing LO P1, P3, A1, A2, C3

[The following information applies to the questions displayed below.]

Craft Pro Machining produces machine tools for the construction industry. The following details about overhead costs were taken from its company records.

Production Activity Indirect Labor Indirect Materials Other Overhead
Grinding $ 370,000
Polishing $ 110,000
Product modification 600,000
Providing power $ 265,000
System calibration 410,000


Additional information on the drivers for its production activities follows.

Grinding 11,000 machine hours
Polishing 11,000 machine hours
Product modification 2,100 engineering hours
Providing power 21,000 direct labor hours
System calibration 1,000 batches
Job 3175 Job 4286
Number of units 140 units 1,750 units
Machine hours 250 MH 2,500 MH
Engineering hours 26 eng. hours 36 eng. hours
Batches 25 batches 75 batches
Direct labor hours 450 DLH 4,050 DLH

Problem 04-3A Part 5

Required:
5.
If the company uses a plantwide overhead rate based on direct labor hours, what is the overhead cost for each unit of Job 3175? Of Job 4286? (Do not round intermediate calculations. Round "OH Cost per unit" answers to 2 decimal places.)
Answer #1
Production Activity Indirect labor Indirect materials Other Overhead Total
Grinding 370,000 370,000
Polishing 110,000 110,000
Product modification 600,000 600,000
Providing power 265,000 265,000
System calibration 410,000 410,000
Total 1,380,000 110,000 265,000 1,755,000

Total budgeted overheads = $1,755,000

Budgeted direct labor hours = 21,000

Predetermined overhead rate = Total budgeted overheads / Budgeted direct labor hour

= 1,755,000/21,000

= $83.5714285714 per direct labor hour

Overhead allocated to job 3175 = Actual direct labor hour x Predetermined overhead rate

= 450 x 83.5714285714

= $37,607

Overhead allocated to job 4286 = Actual direct labor hour x Predetermined overhead rate

= 4,050 x 83.5714285714

= $338,464

Overhead cost per unit of job 3175 = Overhead allocated to job 3175 / Number of units in job 3175

= 37,607/140

= $268.62

Overhead cost per unit of Job 4286 = Overhead allocated to job 4286/ Number of units in job 4286

= 338,464/1,750

= $193.41

Kindly comment if you need further assistance. Thanks‼!